OpenStudy (anonymous):

if you have 10 pies and someone eats 4 and you get 3 times as much how many pies do you have????

OpenStudy (anonymous):

...Really? This kind of question is being asked? What grade is this? Anyways.. Do 10-4*, and you'll get the answer..

OpenStudy (anonymous):

Oops.. 10-4*3.

OpenStudy (anonymous):

10 - 4 = 6 6^3 = 72

OpenStudy (anonymous):

or is it 18......

OpenStudy (anonymous):

Ya it's 18 sorry.

OpenStudy (anonymous):

It's suppose to be 10 - 4 = 6 6 * 3 = 18

OpenStudy (anonymous):

Yeah, it's 18.